# Position: Beyond Euclidean - Foundation Models Should Embrace Non-Euclidean Geometries

### v1

@@ Abstract
In the era of foundation models and Large Language Models (LLMs), Euclidean space has been the de facto geometric setting for machine learning architectures. However, recent literature has demonstrated that this choice comes with fundamental limitations. At a large scale, real-world data often exhibit inherently non-Euclidean structures, such as multi-way relationships, hierarchies, symmetries, and non-isotropic scaling, in a variety of domains, such as languages, vision, and the natural sciences. It is challenging to effectively capture these structures within the constraints of Euclidean spaces. This position paper argues that moving beyond Euclidean geometry is not merely an optional enhancement but a necessity to maintain the scaling law for the next-generation of foundation models. By adopting these geometries, foundation models could more efficiently leverage the aforementioned structures. Task-aware adaptability that dynamically reconfigures embeddings to match the geometry of downstream applications could further enhance efficiency and expressivity. Our position is supported by a series of theoretical and empirical investigations of prevalent foundation models. Finally, we outline a roadmap for integrating non-Euclidean geometries into foundation models, including strategies for building geometric foundation models via fine-tuning, training from scratch, and hybrid approaches.
Machine Learning, ICML
Machine Learning, ICML
@@ 1 Introduction
Foundation models, such as Large Language Models (LLMs), have emerged as a cornerstone of current AI advancements, due to their ability to generalize across diverse tasks with minimal fine-tuning (Bommasani et al., 2021; Devlin et al., 2018; Brown et al., 2020; Ramesh et al., 2021). To date, Euclidean geometry has remained the default framework for designing such models. This preference is largely driven by the natural compatibility of Euclidean geometry with fundamental neural network operations—such as linear transformations, convolutions, and attention mechanisms—all of which can be efficiently executed using standard linear algebra in Euclidean space.
However, real-world datasets often exhibit implicit non-Euclidean structures. These structures manifest in many forms, such as the hierarchical organization of natural language—including concept taxonomies and entailment relationships (Nickel & Kiela, 2017; Tifrea et al., 2019; Le et al., 2019)—as well as hierarchical relationships among object classes, scenes, and their constituent categories in visual data (Ge et al., 2023; Pal et al., 2024). Furthermore, non-Euclidean characteristics are inherent in biological data, such as protein structures (Villegas-Morcillo et al., 2021) and RNA-seq data (Klimovskaia et al., 2020). Given these intrinsically non-Euclidean characteristics of the training data, combined with the ongoing challenges faced by foundation models—from hallucinations to computational inefficiencies—it becomes crucial to question whether Euclidean geometry remain the default for foundational models.
Position: This position paper argues that the development of non-Euclidean foundation models is essential for effectively representing, modeling, and analyzing complex data structures and relationships in real-world applications. With arguments grounded in both theoretical insights and experimental evidence, we contend that by aligning foundation models—whether visual, linguistic, or scientific—with the intrinsic geometries of their training data, we can improve three critical aspects of these models: representational capabilities, adaptability to diverse geometric structures in training datasets, and scalability.
@@
Figure 1: Manifolds with corresponding graph structures or underlying relationships, which represent different types of token relationships: hierarchical (left), uniform (middle), and cyclical (right) dependencies.
@@
Representational Capabilities. Euclidean space has limited capabilities to represent complex geometric structures with diverse local properties, resulting in significant distortion when embedding such data in low-dimensional Euclidean spaces (Tenenbaum et al., 2000). In contrast, hyperbolic spaces, with their negative curvature, excel at representing hierarchical and tree-like structures with minimal distortion in low dimensions (Krioukov et al., 2010; Sarkar, 2011). Similarly, spherical geometries, defined by positive curvature, are well-suited for modeling data with bounded structures and angular relationships (Durrant & Leontidis, 2022; Esteves et al., 2017; Trosten et al., 2023).
@@
Adaptability. Incorporating non-Euclidean geometrical operations into foundation models provides substantially enhanced adaptability to the diverse geometric structures in training datasets, particularly in the case of large-scale datasets—as is typical for these models—where heterogeneity is inherent by design. This adaptability improves the models’ flexibility and transferability, as many instances of downstream tasks benefit from acknowledging the geometric structure in the data, such as knowledge graph completion (Bai et al., 2021; Wang et al., 2021), social network analysis (Yang et al., 2022d; Kennedy et al., 2013), multi-label classification, drug discovery (Poleksic, 2023), user preferences recommendation (Chen et al., 2022b; Yang et al., 2022c, b), and code understanding (Tang et al., 2023).
@@
Low Dimensionality and Scalability. Beyond enhancing representation capabilities, adapting foundation models to the non-Euclidean geometry inherent in their training corpora enables lower-dimensional embeddings, reducing computational costs without sacrificing performance. A critical implication lies in the scaling laws of foundation models (Hoffmann et al., 2022), where performance in Euclidean models follows a power-law scaling of the form $L(N)\propto N^{-\alpha}$, with $L$ and $N$ being the loss and the number of parameters respectively. This scaling behavior reveals inherent inefficiencies in how Euclidean geometries handle increasing model complexity and data dimensionality. In contrast, adopting Riemannian manifolds with tailored curvature properties improves scaling by efficiently compressing information (Brehmer et al., 2024; Kiani et al., 2024). This improvement arises from the manifold’s ability to naturally capture long-range dependencies and hierarchical structures (Tifrea et al., 2019), which the Euclidean space struggles to represent effectively. Moreover, product manifolds and mixed-geometry approaches (Gu et al., 2019; Wang et al., 2021) further improve efficiency and performance by allowing different model components to scale according to their optimal geometric properties.
@@
Moreover, we propose a roadmap for integrating non-Euclidean geometries into foundation models. This includes adapting existing Euclidean models to incorporate these principles and developing foundation models from scratch. We also highlight key challenges and outline the steps required to advance this non-Euclidean vision for foundation models.
@@ 2 Background and Preliminaries
In this section, we give an overview of non-Euclidean spaces, particularly focusing on Riemannian manifolds. For more details please see (Lee, 2013) and Appendix A.
@@ 2.1 Non-Euclidean Geometry Foundations
Riemannian Manifolds. A smooth $n$-dimensional manifold $\mathcal{M}$ is a topological space that is locally Euclidean. Each point $x$ is associated with a tangent space $T_{x}\mathcal{M}$, which is an $n$-dimensional vector space that acts as a first-order local approximation of $\mathcal{M}$. A Riemannian metric $\mathfrak{g}$ on $\mathcal{M}$ is a collection $\mathfrak{g}\coloneqq(\mathfrak{g}_{x})_{x\in\mathcal{M}}$ of positive definite bilinear forms $\mathfrak{g}_{x}(\cdot,\cdot):T_{x}\mathcal{M}\times T_{x}\mathcal{M}\to\mathbb{R}^{n}$, varying smoothly with $x$. $\mathfrak{g}_{x}$ induces the (sectional) curvature at point $x$, which measures how $\mathcal{M}$ deviates from flatness at $x$. A Riemannian manifold is a pairing $(\mathcal{M},\mathfrak{g})$. For example, $\mathbb{R}^{n}$ with the usual Euclidean inner product is a Riemannian manifold with constant curvature 00. $\mathfrak{g}_{x}$ can be seen as a generalization of inner products, where the norm of $p\in T_{x}\mathcal{M}$ is $\|p\|_{\mathfrak{g}}=\sqrt{\mathfrak{g}_{x}(p,p)}$. The choice of $\mathfrak{g}$ induces a global distance function $d(\cdot,\cdot)$ on $\mathcal{M}$. A geodesic between $x,y$ is a local distance minimizing smooth curve. In particular, the shortest paths are geodesics. With certain assumption on the structure of $\mathcal{M}$, one can define the exponential map $\exp_{x}:T_{x}\mathcal{M}\to\mathcal{M}$ for $x\in\mathcal{M}$, and its inverse, the logarithmic map $\log_{x}:\mathcal{M}\to T_{x}\mathcal{M}$. Additionally, the parallel transport map $\mathrm{PT}_{x}(v,w)$, where $v,w\in T_{x}\mathcal{M}$, generalizes translation, transporting $w$ starting at $x$ in the direction of $v$ with no acceleration.
@@ 2.2 Deep Learning in Non-Euclidean Spaces
Recent years have witnessed an increasing interest in extending deep learning techniques to Riemannian manifolds. Here, we discuss several advances for designing neural networks and Transformers in non-Euclidean geometries, as well as optimization on manifolds. A more detailed discussion of recent advances can be found in Appendix A.
Non-Euclidean Neural Networks. Several works have explored neural networks that leverage geodesic distance to perform neural network operations (Bronstein et al., 2017; Masci et al., 2015; Boscaini et al., 2016) and in graph representation (Kong et al., 2022). Within hyperbolic learning, prior works have developed neural network layers (Ganea et al., 2018; Shimizu et al., 2020; Nickel & Kiela, 2017; Chen et al., 2021b; Weber et al., 2020), graph neural networks (Liu et al., 2019a; Chami et al., 2019), vision models (Bdeir et al., 2024; van Spengler et al., 2023), and residual neural networks (He et al., 2025), leveraging more effective embeddings for hierarchical structures. In addition, extensive works have developed equivariant neural networks that encode spherical geometry as inductive bias (Cohen et al., 2018; Esteves et al., 2017; Coors et al., 2018; Defferrard et al., 2020; Esteves et al., 2019). Neural networks for mixed curvature manifolds that encompass both hyperbolic and spherical models have also been proposed (Gu et al., 2019; Bachmann et al., 2020). Many Euclidean convex and stochastic optimization algorithms have been extended to manifold learning as well (Udriste, 1994; Zhang et al., 2016; Bécigneul & Ganea, 2018; Weber & Sra, 2021, 2022).
Non-Euclidean Transformers. Significant advancements have been made toward Transformers in non-Euclidean spaces in recent studies. Prior works have developed attention mechanisms and additional essential operations, such as layer normalization, to develop Transformers in hyperbolic, spherical, and mixed curvature manifolds (Gulcehre et al., 2019; Chen et al., 2021b; Shimizu et al., 2020; Yang et al., 2024c; Lai et al., 2023; Cho et al., 2023).
@@ 3 Foundation Models Should Embrace Non-Euclidean Geometries
Euclidean Foundation Models. Foundation models are typically trained on massive datasets to learn transferable representations that serve as a basis for downstream tasks (Bommasani et al., 2021). Transformer-based language models such as BERT (Devlin et al., 2018), GPT (Brown et al., 2020), and T5 (Raffel et al., 2020) are pre-trained on extensive text corpora. In computer vision, large-scale models such as Vision Transformer (ViT) (Dosovitskiy et al., 2020) and ResNet variants (He et al., 2016) enable state-of-the-art performance on tasks such as image classification and object detection. Multimodal foundation models like CLIP (Radford et al., 2021) and DALL-E (Ramesh et al., 2021) learn cross-domain embeddings for zero-shot image classification and text-to-image generation.
@@ 3.1 Limitations of Euclidean Geometry for Foundation Models
The Euclidean assumption is based on the idea that relationships between data points can be meaningfully characterized using distances measured in a flat, Euclidean space. However, theoretical and experimental verification has demonstrated that Euclidean geometry, with its isotropic nature and uniform scaling, fails to capture the complex structures of real-world data, resulting in significant distortions (Bourgain, 1986; Matoušek, 1996; Arias-de Reyna & Rodríguez-Piazza, 1992; Gupta, 1999; Matoušek, 1999). As a result, high-quality, low-distortion embeddings are often only possible in high-dimensional Euclidean space. Specifically, embeddings of complex structured data, such as hierarchies or trees, provably incur high rates of distortion (Bourgain, 1985; Matoušek, 1996). In this section, we highlight how the flat nature of the Euclidean space results in limitations and challenges for foundational models.
Non-Applicability of the Nash Embedding Theorem. The Nash Embedding Theorem states that any Riemannian manifold $\mathcal{M}$ of dimension $n$ admits an isometric embedding $f$ into $\mathbb{R}^{2n+1}$ (Nash, 1954), seemingly to imply that non-Euclidean spaces would only reduce the embedding dimension by half. However, the isometric embedding here is defined to preserve the Riemannian metric, meaning that it is locally distance preserving—the length of any path is preserved. However, for the shortest path between points $x,y\in\mathcal{M}$, its image under $f$ is not necessarily the shortest path (i.e., Euclidean straight line) between $f(x)$ and $f(y)$. Conversely, measuring the embedding distortion is concerned with whether a map is globally distance preserving. This occurs when the shortest path between $x$ and $y$ remains the shortest path between $f(x)$ and $f(y)$, which is defined by isometric embeddings between metric spaces. Note that an isometric embedding between Riemannian manifolds is in general not an isometric embedding between metric spaces.
In the context of foundational models, we are concerned with global distance-preserving embeddings, as computing the distance between any pairs of token embeddings is crucial for model training. As a result, the Nash Embedding Theorem is not applicable since global distortion could still arise from isometric embeddings between Riemannian manifolds. For this reason, by “isometry”, we refer to those between metric spaces, which captures the (bi-Lipschitz) distortion critical for foundational model embeddings. For more details on distortion and isometric embeddings, see Section A.2.
Dimensionality. High dimensionality is often required to embed complex structures in Euclidean space with low distortion. The following theorem demonstrates the distortion-dimension trade-off that Euclidean embeddings face even in the simple case of unweighted token relationships / self-attention, in the form of complete graphs.
@@
(Matoušek (2002)) Let $X$ be an $n$-point metric space with uniform distance $1$, i.e., an unweighted complete graph with $n$ nodes. For $\epsilon>0$, the minimal $d$ such that $X$ can be embedded into $\mathbb{R}^{d}$ with distortion $(1+\epsilon)$ is $d=\Omega\left(\frac{\log(n)}{\epsilon^{2}\log(1/\epsilon)}\right)$
For any $p<2$, $\epsilon^{2}\log(1/\epsilon)$ tends to 00 at a faster rate than $\epsilon^{p}$ as $\epsilon\to 0$. As a result, Theorem 3.1 implies that $d$ grows near-quadratically with respect to inverse distortion, even when embedding simple structures. Furthermore, note that any un-weighted graph with $n$ nodes can be isometrically embedded into an unweighted complete graph with $n$ nodes. Thus Theorem 3.1 implies the same dimensionality issue for embedding any unweighted graph in Euclidean space.
Distortion. In the cases of more complex structures, non-trivial distortion could even exist regardless of the dimension of the Euclidean space due to its uniform scaling properties. The following theorem implies that a wide range of spaces cannot be isometrically embedded into Euclidean space, based on Markov convexity (Section A.3).
@@
(Lee et al., 2007) Let $(X,d_{X}),(Y,d_{Y})$ be metric spaces. For every $p\in\mathbb{N}$, denote $\Pi_{p}(X),\Pi_{p}(Y)$ the Markov $p$-convexity constant of $X$ and $Y$ respectively. Let $c_{Y}(X)=\inf\{\mathrm{dist}(f):f:X\to Y\}$ denote the minimum distortion of embedding $X$ in $Y$. Then $c_{Y}(X)\geq\frac{\Pi_{p}(X)}{\Pi_{p}(Y)}$.
For instance, when $X$ models hierarchical token relationships, e.g., $X=B_{2^{k}}$ is a complete binary tree of depth $2^{k}$, we have $\Pi_{p}(X)\geq\sqrt{k}$ (Lee et al., 2007). Thus, the distortion for embedding binary trees of depth in any Euclidean space is at least $k$$c_{\mathbb{R}^{n}}(B_{k})\geq\Omega(1)\cdot\sqrt{\log k}$. Similarly, when $X$ represents circular or periodic dependencies in the tokens, e.g., $X$ is a ball of radius $r$ in a vertex-transitive graph, such as a cycle graph, the minimal distortion of embedding $X$ into $\mathbb{R}^{n}$ for any $n$ is $\Omega(\sqrt{\log r})$ (Lee et al., 2007).

### v2

# Position: Beyond Euclidean – Foundation Models Should Embrace Non-Euclidean Geometries

@@ Abstract
In the era of foundation models and Large Language Models (LLMs), Euclidean space has been the de facto geometric setting for machine learning architectures. However, recent literature has demonstrated that this choice comes with fundamental limitations. At a large scale, real-world data often exhibits inherently non-Euclidean structures, such as multi-way relationships, hierarchies, symmetries, and non-isotropic scaling, in a variety of domains, such as languages, vision, and the natural sciences. It is challenging to effectively capture these structures within the constraints of Euclidean spaces. This position paper argues that moving beyond Euclidean geometry is not merely an optional enhancement but a necessity to maintain the scaling law for the next-generation of foundation models. By adopting these geometries, foundation models could more efficiently leverage the aforementioned structures. Task-aware adaptability that dynamically reconfigures embeddings to match the geometry of downstream applications could further enhance efficiency and expressivity. Our position is supported by a series of theoretical and empirical investigations of prevalent foundation models. Finally, we outline a roadmap for integrating non-Euclidean geometries into foundation models, including strategies for building geometric foundation models via fine-tuning, training from scratch, and hybrid approaches.
@@ 1 Introduction
Figure 1: Manifolds and their corresponding graph structures or underlying relationships, which represent different types of token relationships: hierarchical (left), uniform (middle), and cyclical (right) dependencies.
Foundation models, such as Large Language Models (LLMs), have emerged as a cornerstone of current AI advancements due to their ability to generalize across diverse tasks with minimal fine-tuning [bommasani2021on, devlin2018bert, brown2020language, ramesh2021DALLE]. Euclidean geometry has been the default framework for designing such models, largely driven by the natural compatibility of Euclidean geometry with fundamental neural network operations—such as linear transformations, convolutions, and attention mechanisms—which can be executed efficiently using standard linear algebra in Euclidean space. While efficient and intuitive, this choice of Euclidean geometry introduces biases and distortions as a result of its flat geometry and linear assumption, such as in the distances computed for downstream tasks. However, real-world datasets often exhibit implicit non-Euclidean structures, such as the hierarchical organization of natural language—including concept taxonomies and entailment relationships [nickel2017poincare, tifrea2019poincare, le2019inferring]—as well as hierarchical relationships among object classes, scenes, and their constituent categories in visual data [ge2023hyperbolic, pal2024compositional]. Furthermore, non-Euclidean characteristics are inherent in biological data, such as protein structures [villegas2021foldhsphere] and RNA-seq data [klimovskaia2020poincare]. Given the non-Euclidean characteristics of training data, along with the challenges faced by current foundation models—from hallucinations to computational inefficiencies—it becomes crucial to question whether Euclidean geometry should remain the default for foundation models. Existing works that attempt to incorporate non-Euclidean geometry into neural networks focus on finding geometry alternatives to fundamental Euclidean neural network operations that achieves similar goals as the Euclidean counterparts, such as defining neural network layers in non-Euclidean embedding spaces [chen2021fully, yang2024hypformer, HNN++, he2025lresnet] or preserves certain geometric properties [lai2023spherical, cohen2018spherical, deepsphere2020deepsphere] or designing loss functions in a non-Euclidean latent space [ge2023hyperbolic, desai2023hyperbolic]. However, the majority of these works do not actively attempt to address the problem at the level of foundation models and instead focus on low-dimensional settings. See Appendix˜A for further discussion of non-Euclidean geometry and the geometric assumptions of embedding spaces.
Position: The development of non-Euclidean foundation models is essential for effectively representing, modeling, and analyzing complex data structures and relationships in real-world applications. Particularly, this paper advocates for the development of non-Euclidean foundation models at the scale of billions of parameters, which is on both a much larger scale and a much broader geometric scope than existing research that focuses almost entirely on low-dimensional settings within specific geometries, such as the hyperbolic space. Such research requires significant community efforts. Beyond proposing non-Euclidean counterparts of Euclidean operations, which is the current focus, we advocate building the full training infrastructure for this scale, developing and training large-scale non-Euclidean architectures, and analyzing their behavior, which often diverges from low-dimensional models. With arguments grounded in theoretical insights and experimental evidence, we contend that by aligning foundation models—whether visual, linguistic, or scientific—with the intrinsic geometries of their training data, we can improve three critical aspects of these models: representational capabilities, adaptability to diverse geometric structures, and scalability.
@@
Representational Capabilities. Euclidean space has limited capabilities to represent complex geometric structures with diverse local properties, resulting in significant distortion when embedding such data in low-dimensional Euclidean spaces [tenenbaum2000isomap]. In contrast, hyperbolic spaces, with their negative curvature, excel at representing hierarchical structures with minimal distortion in low dimensions [krioukov2010hyperbolic, sarkar2011low]. Similarly, spherical geometries, defined by positive curvature, are well-suited for modeling data with bounded structures and angular relationships [durrant2022hyperspherically, esteves2017learning, trosten2023hubs].
@@
Adaptability. Incorporating non-Euclidean geometrical operations into foundation models provides substantially enhanced adaptability to the diverse geometric structures in training datasets, particularly in the case of large-scale datasets—as is typical for these models—where heterogeneity is inherent by design. This adaptability improves the models’ flexibility and transferability, as many instances of downstream tasks benefit from acknowledging the geometric structure in the data, such as knowledge graph completion [bai2021modeling, wang2021mixed], social network analysis [yang2022hyperbolic, kennedy2013hyperbolicity], multi-label classification, drug discovery [poleksic2023hyperbolic], user preferences recommendation [chen2022modeling, yang2022hrcf, yang2022hicf], and code understanding [tang2023hyperbolic].
@@
Scalability. Adapting foundation models to non-Euclidean geometry enables expressive lower-dimensional embeddings, reducing computational costs without sacrificing performance. A critical implication lies in the scaling laws of foundation models [10.5555/3600270.3602446], where performance in Euclidean models follows a power-law scaling of the form $L(N)\propto N^{-\alpha}$, with $L$ and $N$ being the loss and parameter count. This behavior reveals inherent inefficiencies in how Euclidean space handles increasing model complexity and data dimensionality. In contrast, Riemannian methods have shown promises to improve scaling by efficiently compressing information [brehmer2024does, kiani2024hardness]. For instance, hyperbolic spaces better captures long-range dependencies [tifrea2019poincare] and mixed-curvature approaches [gu2019mixedcurvature, wang2021mixed] allow different model components to scale according to their optimal geometric properties.
@@
Roadmap. Moreover, we propose a roadmap for integrating non-Euclidean geometries into foundation models. This includes both adapting existing Euclidean models to incorporate these principles and developing foundation models from scratch. We also highlight key challenges and outline the steps required to advance this non-Euclidean vision, from architectural design to the creation of non-Euclidean libraries, given that existing frameworks such as DeepSpeed [rasley2020deepspeed] and Flash Attention [dao2022flashattention] are tailored exclusively for Euclidean models.
@@ 2 Background and Preliminaries
In this section, we give an overview of non-Euclidean spaces, particularly focusing on Riemannian manifolds. For more details please see [lee2013manifold] and Appendix˜A.
@@ 2.1 Non-Euclidean Geometry Foundations
Riemannian Manifolds. A smooth $n$-dimensional manifold $\mathcal{M}$ is a topological space that is locally Euclidean. Each point $x$ is associated with a tangent space $T_{x}\mathcal{M}$, which is an $n$-dimensional vector space that acts as a first-order local approximation of $\mathcal{M}$. A Riemannian metric $\mathfrak{g}$ on $\mathcal{M}$ is a collection $\mathfrak{g}\coloneqq(\mathfrak{g}_{x})_{x\in\mathcal{M}}$ of positive definite bilinear forms $\mathfrak{g}_{x}(\cdot,\cdot):T_{x}\mathcal{M}\times T_{x}\mathcal{M}\to\mathbb{R}^{n}$, varying smoothly with $x$. $\mathfrak{g}_{x}$ induces the (sectional) curvature at point $x$, which measures how $\mathcal{M}$ deviates from flatness at $x$. A Riemannian manifold is a pairing $(\mathcal{M},\mathfrak{g})$. For example, $\mathbb{R}^{n}$ with the usual Euclidean inner product is a Riemannian manifold with constant curvature 0. $\mathfrak{g}_{x}$ can be seen as a generalization of inner products, where the norm of $p\in T_{x}\mathcal{M}$ is $\|p\|_{\mathfrak{g}}=\sqrt{\mathfrak{g}_{x}(p,p)}$. The choice of $\mathfrak{g}$ induces a global distance function $d(\cdot,\cdot)$ on $\mathcal{M}$. A geodesic between $x,y$ is a local distance minimizing smooth curve. In particular, the shortest paths are geodesics. With certain assumption on the structure of $\mathcal{M}$, one can define the exponential map $\exp_{x}:T_{x}\mathcal{M}\to\mathcal{M}$ for $x\in\mathcal{M}$, and its inverse, the logarithmic map $\log_{x}:\mathcal{M}\to T_{x}\mathcal{M}$. Additionally, the parallel transport map $\mathrm{PT}_{x}(v,w)$, where $v,w\in T_{x}\mathcal{M}$, generalizes translation, transporting $w$ starting at $x$ in the direction of $v$ with no acceleration.
@@ 2.2 Deep Learning in Non-Euclidean Spaces
Recent years have witnessed an increasing interest in extending deep learning techniques to Riemannian manifolds. We discuss several advances for designing neural networks and Transformers in non-Euclidean geometries, as well as optimization on manifolds, with more details in Appendix˜A.
Non-Euclidean Neural Networks. Several works have explored neural networks that leverage geodesic distance to perform neural network operations [bronstein2017geometric, masci2015geodesic, boscaini2016learning, kong2022geodesic]. Within hyperbolic learning, prior works have developed neural network layers [HNN, HNN++, nickel2017poincare, chen2021fully, weber2020robust], graph neural networks [liu2019HGNN, hgcn2019], vision models [Bdeir2024fully, van2023poincar], and residual neural networks [he2025lresnet]. In addition, extensive works have developed equivariant neural networks that encode spherical geometry as inductive bias [cohen2018spherical, esteves2017learning, coors2018spherenet, deepsphere2020deepsphere, esteves2019equaivariant]. Neural networks for mixed curvature manifolds that encompass both hyperbolic and spherical models have also been proposed [gu2019mixedcurvature, bachmann2020constant]. Many Euclidean convex and stochastic optimization algorithms have been extended to manifold learning as well [udriste1994convex, zhang2016riemannian, becigneul2018riemannian, weber2021projection, weber2022riemannian].
Non-Euclidean Transformers. Significant advancements have been made toward Transformers in non-Euclidean spaces in recent studies. Prior works have developed attention mechanisms and additional essential operations, such as layer normalization, to develop Transformers in hyperbolic, spherical, and mixed curvature manifolds [gulcehre2019hyperbolicAT, chen2021fully, HNN++, yang2024hypformer, lai2023spherical, cho2023curve].
Nevertheless, there is a lack of works for non-Euclidean foundation models. While prior works have attempted or argued for incorporating geometric or topological information into certain aspects of model design [papillon2025beyondeuclid, bronstein2017geometric, hajij2022topological], these prior works almost all focus on low-dimensional settings, with few works that consider pre-trained models [chen2024hypbert], thus omitting the vast potential benefits to be gained from non-Euclidean foundation models.
@@ 3 Foundation Models Should Embrace Non-Euclidean Geometries
Euclidean Foundation Models. Foundation models are typically trained on massive corpora to learn transferable representations that serve as a basis for downstream tasks [bommasani2021on]. Transformer-based language models [devlin2018bert, brown2020language, brown2020language, 2020t5, dubey2024llama3], large-scale vision models such as Vision Transformer (ViT) and ResNet [dosovitskiy2020image, he2016deep], and multimodal foundation models like CLIP [radford2021CLIP] and DALL-E [ramesh2021DALLE], have achieve state-of-the-art performances in a vast amount of tasks across numerous domains.
@@ 3.1 Limitations of Euclidean Geometry for Foundation Models
The Euclidean assumption is that relationships between data points can be meaningfully characterized using distances measured in a flat space. However, theoretical and experimental works have demonstrated that Euclidean geometry, with its isotropic nature and uniform scaling, fails to capture the complex structures of real-world data, resulting in significant distortions [bourgain1986metrical, matouvsek1996distortion, arias1992finite, gupta1999embedding, matouvsek1999embedding]. As a result, high-quality, low-distortion embeddings are often only possible in high-dimensional Euclidean space. Specifically, embeddings of complex structured data, such as hierarchies or trees, provably incur high rates of distortion [bourgain1985lipschitz, matouvsek1996distortion]. In this section, we highlight how the flat nature of the Euclidean space results in limitations and challenges for foundational models.
Non-Applicability of the Nash Embedding Theorem. The Nash Embedding Theorem states that any Riemannian manifold $\mathcal{M}$ of dimension $n$ admits an isometric embedding $f$ into $\mathbb{R}^{2n+1}$ [nash1954embedding], seemingly to imply that non-Euclidean spaces would only reduce the embedding dimension by half. However, the isometric embedding here is defined to preserve the Riemannian metric, meaning that it is locally distance preserving—the length of any path is preserved. However, for the shortest path between points $x,y\in\mathcal{M}$, its image under $f$ is not necessarily the shortest path (i.e., Euclidean straight line) between $f(x)$ and $f(y)$. Conversely, measuring the embedding distortion is concerned with whether a map is globally distance preserving w.r.t the ambient space, or when the shortest path between $x$ and $y$ remains the shortest path between $f(x)$ and $f(y)$, which is defined by isometric embeddings between metric (ambient) spaces. Note that an isometric embedding between Riemannian manifolds is in general not an isometric embedding between metric spaces.
We are concerned with global distance-preserving embeddings for foundational models, as the distance between any pair of token embeddings is crucial for model training. Thus, the Nash Embedding Theorem is not applicable since global distortion could still arise from isometric embeddings between Riemannian manifolds. For this reason, by “isometry”, we refer to those between metric spaces. See Section˜A.2 for more details. As the Nash Embedding Theorem is not applicable, Euclidean embeddings suffer from several limitations, which we detail below.
Dimensionality. Euclidean space requires high dimensionality to embed complex structures with low distortion, which ties directly to high demand for model size and computational resources. The following theorem shows the distortion-dimension tradeoff for Euclidean embeddings even in the simple case of unweighted token relationships, in the form of complete graphs.
@@
(matousek2002lectures) Let $X$ be an $n$-point metric space with uniform distance $1$, i.e., an unweighted complete graph with $n$ nodes. For $\epsilon>0$, the minimal $d$ such that $X$ can be embedded into $\mathbb{R}^{d}$ with distortion $(1+\epsilon)$ is $d=\Omega\left(\frac{\log(n)}{\epsilon^{2}\log(1/\epsilon)}\right)$
For any $p<2$, $\epsilon^{2}\log(1/\epsilon)$ tends to 0 faster than $\epsilon^{p}$ as $\epsilon\to 0$. As a result, Theorem˜3.1 implies that $d$ grows near-quadratically w.r.t. inverse distortion. Furthermore, any unweighted graph with $n$ nodes can be isometrically embedded into an unweighted complete graph with $n$ nodes. Thus Theorem˜3.1 implies the same dimensionality issue for embedding any unweighted graph in Euclidean space.
Distortion. Non-trivial distortion could exist regardless of the dimension of the Euclidean space in the cases of more complex structures. The following theorem implies that a wide range of spaces cannot be isometrically embedded into Euclidean space, based on Markov convexity (Section˜A.3).
@@
[lee2009trees] Let $(X,d_{X}),(Y,d_{Y})$ be metric spaces. For every $p\in\mathbb{N}$, denote $\Pi_{p}(X),\Pi_{p}(Y)$ the Markov $p$-convexity constant of $X$ and $Y$ respectively. Let $c_{Y}(X)=\inf\{\mathrm{dist}(f):f:X\to Y\}$ denote the minimum distortion of embedding $X$ in $Y$. Then $c_{Y}(X)\geq\frac{\Pi_{p}(X)}{\Pi_{p}(Y)}$.
When $X$ models hierarchical token relationships, e.g., $X=B_{2^{k}}$ is a complete binary tree of depth $2^{k}$, the distortion for embedding binary trees of depth in any Euclidean space is at least $\Omega(1)\cdot\sqrt{\log k}$. When $X$ represents circular or periodic dependencies in tokens, e.g., $X$ is a ball of radius $r$ in a vertex-transitive graph, the minimal distortion of embedding $X$ into $\mathbb{R}^{n}$ for any $n$ is $\Omega(\sqrt{\log r})$ [lee2009trees].